# GRPO-MA: Multi-Answer Generation in GRPO for Stable and Efficient Chain-of-Thought Training

### v2

@@ Abstract
Recent progress, such as DeepSeek-R1, has shown that the GRPO algorithm, a Reinforcement Learning (RL) approach, can effectively train Chain-of-Thought (CoT) reasoning in Large Language Models (LLMs) and Vision-Language Models (VLMs). In this paper, we analyze three challenges of GRPO: gradient coupling between thoughts and answers, sparse reward signals caused by limited parallel sampling, and unstable advantage estimation. To mitigate these challenges, we propose GRPO-MA, a simple yet theoretically grounded method that leverages multi-answer generation from each thought process, enabling more robust and efficient optimization. Theoretically, we show that the variance of thought advantage decreases as the number of answers per thought increases. Empirically, our gradient analysis confirms this effect, showing that GRPO-MA reduces gradient spikes compared to GRPO. Experiments on math, code, and diverse multimodal tasks demonstrate that GRPO-MA substantially improves performance and training efficiency. Our ablation studies further reveal that increasing the number of answers per thought consistently enhances model performance.
@@ 1 Introduction
The Deepseek-R1 model demonstrates that reinforcement learning (RL)–specifically through Group Relative Policy Optimization (GRPO) Shao et al. (2024)—is effective in training Chain-of-Thought (CoT). This method involves prompting the Large Language Model (LLM) to generate a reasoning trace before producing a final answer, which is then reinforced via a reward signal to enhance the model’s reasoning capabilities. Subsequently, methods such as DAPO Yu et al. (2025), Dr.GRPO Liu et al. (2025), and GPG Chu et al. (2025) have improved upon GRPO’s loss function from various perspectives, achieving more stable training curves and better results on mathematical problems. Beyond textual reasoning tasks, the GRPO paradigm has also been extended to multi-modal scenarios Chen et al. (2025b); Shen et al. (2025); Huang et al. (2025b); Feng et al. (2025); Song et al. (2025); Kim et al. (2025). Compared to refinements of the algorithm itself, these multi-modal applications have largely adopted task-specific reward functions to improve performance on specific objectives, such as adding a temporal video reward in Video-R1 Feng et al. (2025) or a trajectory distance reward in ManipVLM-R1 Song et al. (2025). These studies have shown that CoT, trained with verifiable reward functions in RL, significantly enhances multi-modal reasoning. Despite these successes, GRPO still suffers from inherent challenges that limit training stability, efficiency, and overall effectiveness. These challenges manifest in several distinct yet interrelated aspects of the GRPO framework—namely, gradient coupling between thoughts and answers, inefficiencies in generation, and instability in advantage estimation.
A well-known issue is the mismatch between reasoning traces and final answers: the reasoning may be valid while the final answer is wrong, or conversely, a flawed reasoning may still yield a correct answer. This phenomenon can be observed in both pure textual reasoning tasks Simoni et al. (2025); Lin et al. (2025a); Paul et al. (2024); Turpin et al. (2023) and multi-modal tasks Chen et al. (2025b); Balasubramanian et al. (2025) including our experiments 5.4. Since the gradients of thoughts and answers are inherently coupled in GRPO, such inconsistencies can distort the gradient direction and consequently undermine training effectiveness. Although GRPO-CARE Chen et al. (2025b) introduces a consistency reward to alleviate this, it risks reward hacking and is difficult to apply when semantic consistency is ill-defined (e.g., it is difficult to judge the consistency between a CoT and the numerical coordinates of a predicted bounding box).
The second challenge arises from the high computational cost of response generation in GRPO. For each problem instance, the algorithm must sample multiple complete CoT–answer pairs, which becomes prohibitive when the task is difficult or the answer space is large. In such settings, the limited number of samples often fails to capture the rare successful outcomes, yielding groups where all rewards are zero. Since GRPO computes advantages by normalizing rewards within each group, an all-zero reward set collapses into uniform zero advantages, providing no relative signal. This collapse of advantage values eliminates informative gradients and stalls optimization. A related approach, DAPO, attempts to address this issue by continually sampling until at least one positive reward is found, thereby avoiding advantage collapse. However, this strategy incurs extremely high computational overhead when rewards are sparse. Hence, an efficient mechanism is needed to enrich reward signals without incurring excessive generation cost.
The third challenge concerns variance in advantage estimation. From a probabilistic modeling perspective, a “good” thought is fundamentally one that increases the likelihood of generating a “good” answer. It follows that the most robust method for evaluating a thought’s quality would be to assess the overall distribution of multiple answers sampled from it. The GRPO, in its current design, estimates a thought’s advantage based on a single sampled answer. A consequence of this single-sample estimation, particularly when combined with the stochasticity of high-temperature sampling in LLMs and VLMs, is the potential for increased variance in the advantage estimation. Crucially, more accurate estimation of thought advantages is not only beneficial for reducing training instability, but also for guiding the model to internalize what constitutes a genuinely “good” thought. This, in turn, enables the model to generate higher-quality answers more reliably.
In this paper, we propose GRPO-MA (GRPO with Multi-Answer). For each of $K$ thoughts, we sample $M$ answers. A thought’s value is the average reward of its $M$ answers, which is used to derive its advantage relative to other thoughts. while each of the $K\times M$ answers also receives its own advantage. These two advantages are then used to update thought and answer tokens separately. Our theoretical analysis, based on the delta method Oehlert (1992), reveals the distinct effects of $K$ and $M$ on the variance of the thought’s advantage. The analysis shows that as $M$ increases, the variance monotonically decreases towards zero. In contrast, increasing $K$ only reduces the variance to a non-zero constant. This design brings three benefits: (1) It reduces gradient coupling from noisy thought–answer mismatches by basing thought updates on an averaged reward. (2) The multi-answer estimate of a thought’s value has lower variance, leading to more stable advantage estimation. (3) It is computationally efficient by amortizing the cost of generating $K$ thoughts across $M$ answers each, avoiding the higher cost of generating $K\times M$ full reasoning responses, while still preserving a diverse and informative set of reward signals.
We evaluate the effectiveness of GRPO-MA on Code, Math, several distinct vision tasks (Object Detection, Affordance Prediction, Trajectory Prediction, Demand Prediction, OCR-based VQA) and a simulator-based visual manipulation task. Compared to a GRPO baseline with $K$ responses, GRPO-MA yields clear gains with only a marginal increase in training time. Compared to a baseline with $K\times M$ responses, it achieves similar or slightly better performance using only about 60% of the training time, highlighting improved sample efficiency from more stable gradient estimation. In the visual manipulation task with extremely sparse rewards, GRPO-MA substantially outperforms the standard GRPO algorithm. Our ablation studies further show that increasing $M$ generally leads to improved model performance, and that the stability gained from more reliable thought advantage estimation may play an even more critical role than the sheer richness of reward signals. Finally, we also compare gradient spikes during training and find that GRPO-MA produces fewer gradient spikes, reflecting greater stability in the training progress.
In summary, our contributions are as follows:
  • We propose the GRPO-MA algorithm, a simple but effective and general improvement strategy for GRPO that is compatible with other mainstream enhancements like DAPO.
  • We provide a theoretical analysis showing that our method can improve the stability of advantage estimation, leading to more stable gradients.
  • Across multiple distinct tasks, GRPO-MA consistently achieves performance gains over the baseline GRPO.
@@ 2 Related Work
The GRPO algorithm has inspired several works to enhance its stability and efficiency by refining its loss function and sampling strategies. DAPO Yu et al. (2025) introduces several “tricks” to stabilize training, such as Clip-Higher for exploration, Dynamic Sampling to filter uninformative samples, and a Token-Level Policy Gradient Loss to properly weight complex reasoning chains. Dr. GRPO Liu et al. (2025) corrects inherent response length bias and question difficulty bias by removing specific normalization terms from the loss and advantage calculation, leading to more stable training. Generative Policy Gradient (GPG) Chu et al. (2025) simplifies the GRPO objective and introduces a gradient rescaling method to counteract “zero-gradient” samples, ensuring more effective policy updates. Further research has focused on improving efficiency, with CPPO Lin et al. (2025b) pruning low-impact samples to reduce computational cost and Off-Policy GRPO Mroueh et al. (2025) using stale data to improve sample efficiency. Other works enhance stability, such as GSPO Zheng et al. (2025), which realigns importance sampling at the sequence level; GMPO Zhao et al. (2025), which uses a geometric mean to mitigate sensitivity to outliers ; and GTPO Simoni et al. (2025), which resolves gradient conflicts and prevents policy collapse through trajectory analysis. Additionally, specialized solutions like Spectral Policy Optimization Chen et al. (2025a) create learning signals for “all-negative” sample groups using AI feedback.
@@ 3 Preliminary: GRPO
The GRPO algorithm Shao et al. (2024) is a Proximal Policy Optimization (PPO) variant Schulman et al. (2017). As a model-free method, GRPO omits a value model and instead calculates advantage values by directly normalizing the rewards obtained from generated responses.
For a given prompt $p$, GRPO samples $K$ responses $O=\{o_{1},\dots,o_{K}\}$ from $\pi_{\theta}$, obtain rewards $\{R_{1},\dots,R_{K}\}$, and compute the advantage as $A(o_{i})=\tfrac{R_{i}-\mathrm{Mean}(\{R_{k}\})}{\mathrm{Std}(\{R_{k}\})}$. In general, a response $o_{i}$ consists of a thought $th_{i}$ and an answer $ans_{i}$.
The GRPO objective is built upon a clipped surrogate objective that maximizes the expected advantage while regularizing the policy change. Formally, given a generated sequence $y$, the clipped objective is defined as:
$\mathcal{J}_{\text{clip}}(\theta,y)=\tfrac{1}{|y|}\sum_{t=1}^{|y|}\min\!\big(r_{t}(\theta)A(y),\operatorname{clip}(r_{t}(\theta),1-\varepsilon,1+\varepsilon)A(y)\big)-\beta D_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})$ (1)
where $r_{t}(\theta)=\tfrac{\pi_{\theta}(y_{t}|p,y_{<t})}{\pi_{\theta_{\text{old}}}(y_{t}|p,y_{<t})}$, $y$ represents a sequence of generated tokens and $\pi_{\text{ref}}$ is a fixed reference model. Building on this definition, the overall GRPO objective aggregates over multiple sampled outputs:
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{\begin{subarray}{c}(p,a)\sim\mathcal{D}\\ \{o_{i}\}_{i=1}^{K}\sim\pi_{\theta_{\text{old}}}(o|p)\end{subarray}}\left[\tfrac{1}{K}\sum_{i=1}^{K}\mathcal{J}_{\text{clip}}(\theta,o_{i})\right]$ (2)
@@ 4.1 Pipeline of GRPO-MA
Figure 1: The operational flow of advantage estimation in GRPO and GRPO-MA. In the baseline GRPO framework (top), the advantage is computed from a single thought–answer pair, inherently coupling the estimation of thought and answer advantages to a single reward signal. In contrast, GRPO-MA (bottom) extends this setting by sampling multiple answers for each thought. This design decouples the estimation of thought and answer advantages and leverages aggregated information from multiple reward signals, thereby yielding richer supervision and enabling more robust and stable estimation of thought-level advantages.
@@
$\mathcal{J}_{\text{GRPO-MA}}(\theta)=\mathbb{E}_{\begin{subarray}{c}(p,a)\sim\mathcal{D},\\ \{th_{i}\}_{i=1}^{K}\sim\pi_{\theta_{\text{old}}}(th\mid p),\\ \{ans_{i,j}\}_{i=1}^{K}{}_{j=1}^{M}\sim\pi_{\theta_{\text{old}}}(ans\mid p,th)\end{subarray}}\left[\tfrac{1}{K}\sum_{i=1}^{K}\mathcal{J}_{\text{clip}}(\theta,th_{i})+\tfrac{1}{KM}\sum_{i=1}^{K}\sum_{j=1}^{M}\mathcal{J}_{\text{clip}}(\theta,ans_{i,j})\right]$ (3)
where the first term updates the policy based on the thoughts, using the thought advantage $A(th_{i}$) within its $\mathcal{J}_{\text{clip}}$ calculation, and the second term updates the policy based on the answers, using the answer advantage $A(ans_{i,j})$.
@@ 4.2.1 Preliminaries
Fix a prompt $p$. Thoughts $th_{1},\dots,th_{K}$ are sampled independently from a distribution $\pi_{\theta}(\cdot\mid p)$ conditioned on the prompt $p$.
For each thought $th_{i}$ we generate $M$ answers independently from the conditional policy, written as $\text{ans}_{i,j}\overset{\text{i.i.d.}}{\sim}\pi_{\theta}(\cdot\mid p,th_{i}),\ j=1,\dots,M$. Here the index $j$ simply labels different samples, all drawn from the same distribution $\pi(\cdot\mid p,th_{i})$.
Each answer $\text{ans}_{i,j}$ is evaluated by a reward function $r$. We denote the resulting reward as a random variable $R_{i,j}:=r(\text{ans}_{i,j},p)$. $R_{i,j}$ is random because the answer $\text{ans}_{i,j}$ itself is sampled.
Conditioned on a given thought $th_{i}$ and prompt $p$, the rewards $R_{i,j}$ are i.i.d. with mean $\mu_{R_{i}}$ and variance $\sigma_{R_{i}}^{2}$. The empirical value estimator of thought $th_{i}$ is the sample mean $V(th_{i})=\tfrac{1}{M}\sum_{j=1}^{M}R_{i,j}$, with $\mathbb{E}[V(th_{i})]=\mu_{R_{i}}$ and $\mathrm{Var}(V(th_{i}))=\tfrac{\sigma_{R_{i}}^{2}}{M}$.
We assume the thought-value estimates $V(th_{1}),\dots,V(th_{K})$ are independent, i.e., the covariance matrix is diagonal. In practice, small correlations may exist since all thoughts are sampled from the same prompt. However, Appendix A.2.5 shows that the diagonal entries capture most of the covariance energy, suggesting this assumption is largely reasonable.
Finally, define the sample mean and standard deviation across thoughts as $\bar{V}=\tfrac{1}{K}\sum_{k=1}^{K}V(th_{k})$, $S_{V}=\sqrt{\tfrac{1}{K-1}\sum_{k=1}^{K}(V(th_{k})-\bar{V})^{2}}$, and the thought advantage as $A(th_{i})=\tfrac{V(th_{i})-\bar{V}}{S_{V}}$.
4.2.2 Variance of the Thought Advantage

### v3

# Why Tree-Style Branching Matters for Thought Advantage Estimation in GRPO

@@ Abstract
Group Relative Policy Optimization (GRPO) trains Chain-of-Thought reasoning with verifiable rewards, but estimating thought-level advantages without value functions often suffers from high variance. Although tree-style branching is used in practice to reduce the variance, it lacks a theoretical explanation of why it works and whether it is important or even potentially necessary. We study thought-level advantage estimation in GRPO from a variance perspective under a minimal tree-style setting where multiple answers are sampled for each thought. Using the multivariate delta method, we reveal an asymmetry in how different sampling dimensions affect variance. Increasing the number of sampled thoughts ($K$) leaves a strictly positive variance floor, whereas increasing the number of answers per thought ($M$) induces a monotonic decrease in variance, asymptotically decreasing it to zero. This implies that accurate thought-level advantage estimation is impossible through scaling thought sampling alone, making branching a potentially necessary mechanism rather than a heuristic. Experiments further provide empirical evidence for both the effectiveness and necessity of answer-level branching, demonstrating improved optimization stability, training efficiency, and final performance not only in math but also across a broad range of vision domains and under different model architectures and sizes.
@@ 1 Introduction
The DeepSeek-R1 model demonstrates that reinforcement learning (RL)—particularly Group Relative Policy Optimization (GRPO) (Shao et al., 2024)—is effective for training Chain-of-Thought (CoT) reasoning. GRPO prompts the LLM to generate a reasoning trace before producing the final answer and then reinforces this process via verifiable rewards. Subsequent methods such as DAPO (Yu et al., 2025), Dr.GRPO (Liu et al., 2025), and GPG (Chu et al., 2025) refine GRPO’s loss function from different perspectives, achieving more stable training and stronger mathematical reasoning. Beyond text-based tasks, the GRPO paradigm has also expanded to multi-modal domains (Chen et al., 2025; Shen et al., 2025; Huang et al., 2025b; Feng et al., 2025; Song et al., 2025; Kim et al., 2025), where task-specific reward designs are adopted to improve performance under diverse objectives. Collectively, these results demonstrate that CoT combined with verifiable RL rewards substantially enhances reasoning across modalities.
Although GRPO has demonstrated strong empirical performance across a wide range of domains, it still suffers from instability in thought-level advantage estimation. In standard GRPO-style algorithms, each thought is typically evaluated using a single sampled answer, which contributes to high-variance advantage estimates and unstable gradient updates during training. A growing body of recent work (Kazemnejad et al., 2024; Li et al., 2025; Yang et al., 2025; Guo et al., 2025) empirically adopts tree-style branching to improve the stability of GRPO-style training. However, most of these tree-style methods primarily demonstrate effectiveness through empirical results only on mathematical tasks, rely on complex heuristic segmentation or branching rules, and lack a principled theoretical explanation of why branching improves learning. More importantly, they do not address whether such branching is merely a helpful heuristic or necessary for accurate thought advantage estimation when using GRPO-style, group-normalized advantage estimators without value models.
In this work, we build upon the empirical successes of tree-style branching in GRPO-style training and take a complementary theoretical perspective, focusing on the variance properties of thought-level advantage estimation. To facilitate principled analysis, we simplify the branching mechanism and consider a minimal form of tree search: instead of introducing complex segment- or step-level branching rules, we perform branching only at the answer level. Concretely, we delimit each thought using an explicit XML tag (e.g., </think>) and sample multiple answers conditioned on the same thought. We name this mechanism GRPO-MA (GRPO with Multi-Answer).
Under this formulation, we analyze thought-level advantage estimation in GRPO from a variance perspective and derive an approximate variance characterization using the multivariate delta method. Our analysis goes beyond the observation that increasing the total number of samples reduces variance, and instead reveals a fundamental asymmetry between different sampling dimensions. Concretely, under GRPO-style group-normalized estimators without value models, increasing the number of independently sampled thoughts ($K$) alone cannot eliminate thought advantage estimation noise: the variance converges to a strictly positive constant, regardless of how large $K$ becomes. In contrast, increasing the number of answers per thought ($M$) suppresses the estimation noise along a different sampling dimension, monotonically decreasing the thought-level advantage variance to zero. Together, these results clarify the distinct statistical roles of the two sampling dimensions and explain why tree-style branching is consistently effective in practice. More importantly, they establish that such branching is not merely a heuristic, but a principled and potentially necessary mechanism for accurate thought-level advantage estimation under GRPO-style estimators without value models.
Beyond theory, our experiments using GRPO-MA provide direct empirical evidence for both the effectiveness and the necessity of answer-level branching in GRPO-style training. Comparing T4A1111TKKAMM denotes a GRPO-style sampling configuration with $K$ thoughts and $M$ answers per thought; M=1M{=}1 corresponds to standard GRPO, while M>1M{>}1 corresponds to GRPO-MA. The same notation is used in the experiments. and T4A4, we show that answer-level branching consistently improves performance across diverse domains. More importantly, comparing T4A4 with a thought-scaling baseline T16A1 reveals a strictly better efficiency–performance trade-off: despite operating at substantially lower computational cost, T4A4 achieves slightly better final performance. This indicates that answer-level branching is not merely an efficiency optimization, but a necessary design choice under GRPO-style training, as it enables simultaneous gains in both efficiency and performance. Moreover, T4A4 exhibits significantly fewer gradient spikes than T16A1, as measured by the Gradient Spike Score (GSS), consistent with our variance analysis that lower-variance advantage estimation leads to more stable optimization. Across multiple tasks, model sizes, and architectures, these benefits are robust and not task- or model-specific. Our contributions are summarized as follows:
  • Asymmetric variance in advantage estimation. We present the first variance-based analysis of thought-level advantage estimation in GRPO-style algorithms using the multivariate delta method. Our analysis reveals a fundamental asymmetry in how different sampling dimensions affect the variance of the estimated advantages, explaining why branching is effective and potentially necessary for reliable advantage estimation under GRPO-style estimators.
  • GRPO-MA: a minimal tree-style instantiation. Motivated by this analysis, we propose GRPO-MA, a lightweight and stable instantiation of tree-style branching that introduces answer-level branching via explicit semantic delimiters natively supported by GRPO, requiring minimal modifications and virtually no additional hyperparameters.
  • Broad empirical effectiveness across tasks and model configurations. Extensive experiments across mathematics, code generation, vision, and embodied manipulation, as well as models of different sizes and architectures, demonstrate that GRPO-MA consistently outperforms standard GRPO and achieves a strictly better efficiency–performance trade-off than thought-scaling GRPO baselines..
@@ 2 Related Work
2.1 GRPO Stabilization via Objective and Gradients
A line of work improves the stability and efficiency of GRPO-style training by revisiting the objective, advantage computation, and gradient estimation. DAPO (Yu et al., 2025) combines several practical techniques, including exploration-aware clipping (Clip-Higher), dynamic filtering of uninformative samples (Dynamic Sampling), and a token-level policy-gradient loss to better weight long reasoning traces. Dr. GRPO (Liu et al., 2025) mitigates response-length and question-difficulty biases by removing specific normalization terms in the loss and advantage calculations. GPG (Chu et al., 2025) simplifies the GRPO objective and introduces gradient rescaling to reduce the impact of “zero-gradient” samples. Related methods further target stability from complementary angles: GSPO (Zheng et al., 2025) realigns importance sampling at the sequence level, GMPO (Zhao et al., 2025) uses geometric-mean aggregation to reduce sensitivity to outliers, and GTPO (Simoni et al., 2025) analyzes trajectories to resolve gradient conflicts and prevent policy collapse.
2.2 Tree Search in LLMs Reinforcement Learning
Recent research has shifted from independent chain-based rollouts to tree-structured sampling to enhance exploration efficiency and credit assignment. VinePPO (Kazemnejad et al., 2024) performing Monte Carlo value estimation through multiple suffix rollouts. Evolving from this, SPO (Guo et al., 2025) generalizes the concept of segmentation, enabling flexible granularity advantage estimation. TreeRL (Hou et al., 2025) introduced an entropy-guided search strategy that expands the search at high-uncertainty tokens. TreePO (Li et al., 2025) estimating advantages based on the collective outcomes of grouping sub-trees, improving robustness. TreeRPO (Yang et al., 2025) defines tree nodes at the level of reasoning steps and derive step-wise supervision by propagating outcome rewards across sibling branches.
A key distinction among prior tree-search methods lies in where branching is introduced along the generation process. In this work, we adopt semantic XML tags (e.g., </think>) as branching points, which provide a simple and implementation-friendly instantiation. However, the primary goal of this paper is not to advocate a particular branching heuristic as optimal. Instead, we intentionally focus on a simplified and concrete branching scheme to facilitate a principled analysis of why branching is effective, and more importantly, why it is potentially necessary under the GRPO-style advantage estimator.
@@ 3 Preliminary: GRPO
GRPO (Shao et al., 2024) is a PPO-style algorithm (Schulman et al., 2017) that computes advantages by normalizing rewards from $K$ sampled responses. For a prompt $p$, GRPO generates responses $\{o_{i}\}_{i=1}^{K}$ with rewards $\{R_{i}\}$ and computes $A(o_{i})=\tfrac{R_{i}-\mathrm{Mean}(\{R_{k}\})}{\mathrm{Std}(\{R_{k}\})}$. The complete GRPO objective can be written as:
$\mathcal{J}_{\text{GRPO}}(\theta)$ (1)
$=\;\operatorname*{\mathbb{E}}_{\begin{subarray}{c}(p,a)\sim\mathcal{D}\\ o\sim\pi_{\theta_{\text{old}}}\end{subarray}}\Bigg[\frac{1}{K}\sum_{i=1}^{K}\frac{1}{T^{o}_{i}}\sum_{t=1}^{T^{o}_{i}}\ell_{\text{clip}}\!\big(r_{t},A(o_{i})\big)\Bigg]$
$-\;\beta\,D_{\mathrm{KL}}(\pi_{\theta}\|\pi_{\text{ref}}).$
where $T^{o}_{i}$ denotes the length of the $i$-th output trajectory $o_{i}=(o_{i,1},\dots,o_{i,T^{o}_{i}})$. The term $r_{t}$ is the per-token likelihood ratio, $r_{t}=\frac{\pi_{\theta}(o_{i,t}\mid p,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid p,o_{i,<t})}$, where $\pi_{\theta}$ and $\pi_{\theta_{\text{old}}}$ denote the current and behavior policies used for on-policy sampling. We use the shorthand $\ell_{\text{clip}}(r,A)\triangleq\min\!\big(rA,\operatorname{clip}(r,1{\pm}\varepsilon)A\big)$ for the clipped surrogate term.
@@ 4.1 Pipeline of GRPO-MA
Figure 1: The operational flow of advantage estimation in GRPO and GRPO-MA. In the baseline GRPO framework (top), the advantage is computed from a single thought–answer pair, inherently coupling the estimation of thought and answer advantages to a single reward signal. In contrast, GRPO-MA (bottom) extends this setting by sampling multiple answers for each thought.
@@
$\mathcal{J}_{\text{GRPO-MA}}(\theta)$ (2)
$=\;\operatorname*{\mathbb{E}}_{\begin{subarray}{c}(p,a)\sim\mathcal{D}\\ \mathrm{th}\sim\pi_{\theta_{\text{old}}}\end{subarray}}\Bigg[\frac{1}{K}\sum_{i}\frac{1}{T^{\mathrm{th}}_{i}}\sum_{t=1}^{T^{\mathrm{th}}_{i}}\ell_{\text{clip}}\!\big(r_{t},A(\mathrm{th}_{i})\big)\Bigg]$
$+\;\operatorname*{\mathbb{E}}_{\mathrm{ans}\sim\pi_{\theta_{\text{old}}}}\Bigg[\frac{1}{KM}\sum_{i,j}\frac{1}{T^{\mathrm{ans}}_{i,j}}\sum_{t=1}^{T^{\mathrm{ans}}_{i,j}}\ell_{\text{clip}}\!\big(r_{t},A(\mathrm{ans}_{i,j})\big)\Bigg]$
$-\;\beta\,D_{\mathrm{KL}}(\pi_{\theta}\|\pi_{\text{ref}}).$
where $T^{\mathrm{th}}_{i}$ and $T^{\mathrm{ans}}_{i,j}$ denote the lengths of the thought trajectory $\mathrm{th}_{i}$ and the answer trajectory $\mathrm{ans}_{i,j}$, respectively. The term $A(\mathrm{th}_{i})$ is the thought-level advantage computed by aggregating the rewards of the $M$ answers associated with $\mathrm{th}_{i}$, while $A(\mathrm{ans}_{i,j})$ denotes the advantage of an individual answer. The remaining notation follows the same definitions as in the GRPO objective.
@@ 4.2.1 Preliminaries
Fix a prompt $p$. We sample $K$ thoughts $th_{i}\overset{\text{i.i.d.}}{\sim}\pi_{\theta}(\cdot\mid p)$, and for each $th_{i}$ sample $M$ answers $\text{ans}_{i,j}\sim\pi_{\theta}(\cdot\mid p,th_{i})$. Each answer is scored by a deterministic reward function, defining $R_{i,j}:=r(p,\text{ans}_{i,j})$.
Let $\mu_{R_{i}}:=\mathbb{E}[R_{i,j}\mid p,th_{i}]$ and $\sigma_{R_{i}}^{2}:=\mathrm{Var}(R_{i,j}\mid p,th_{i})$, where $\{R_{i,j}\}_{j=1}^{M}$ are i.i.d. conditional on $(p,th_{i})$.
The empirical thought value $V(th_{i}):=\tfrac{1}{M}\sum_{j=1}^{M}R_{i,j}$ then satisfies $\mathbb{E}[V(th_{i})\mid p,th_{i}]=\mu_{R_{i}}$ and $\mathrm{Var}(V(th_{i})\mid p,th_{i})=\sigma_{R_{i}}^{2}/M$.
Independence assumption.
We assume $\{V(th_{i})\}_{i=1}^{K}$ are mutually independent, i.e., the covariance matrix is diagonal. While mild correlations may arise from sharing the same prompt, Appendix C.5 shows that diagonal terms dominate the covariance structure, making this assumption reasonable.
4.2.2 Approximate Variance Estimation